OpenStudy (anonymous):

an invested a total of $3,000 in two mutual funds.one find earned a 6% profit while the other earned a 2% profit.if the investors total profit was $132,how much was invested in each mutual fund? the amount invested in the mutual fund the earned 6% was $____________. the amount invested in the mutual fund the earned 2% was $____________.

OpenStudy (anonymous):

im lost

OpenStudy (dumbcow):

its solving a system of 2 equations...setting it up can be hard part at first

OpenStudy (mathstudent55):

You need a variable to represent the amount invested at 6%. Let's call that x. x amount was invested at 6%. Since the total amount was 3000, the amount left to be invested at 2% is 3000 - x. Ok so far?

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

Now let's see how much interest is earned by each part of the investment. x invested at 6% earns 6% of x which is 0.06x. (3000 - x) invested at 2% earns 2% of (3000 - x) which is 0.02(3000 - x).

OpenStudy (mathstudent55):

We know the total amount earned in interest is $132, so we can write the equation: 0.06x + 0.02(3000 - x) = 132

OpenStudy (mathstudent55):

Now you need to solve for x, to find the amount invested at 6%. Then subtract x from 3000 to find the amount invested at 2%.

OpenStudy (anonymous):

x=1800

OpenStudy (anonymous):

x=1200

OpenStudy (mathstudent55):

0.06x + 0.02(3000 - x) = 132 0.06x + 60 - 0.02x = 132 0.04x = 72 x = 1800 Correct. Then 3000 - x = 3000 - 1800 = 1200

OpenStudy (mathstudent55):

The amount invested at 6% is $1800. The amount invested at 2% is $1200.